the specific heat of a certain type of metal is 0.128 J/(g C). What is the final temperature if 305 J of heat is added to 39.3 g of this metal, initially at 20.0 C?
Answer #1

Answex briven that Specific heat C=0.12817(9°c) q=305 J mass m = 39.39 initial temperature 1,=20°C Heste, 9= m((T2-11) 305= 3